# Unifying uncertainties for rotor-like quantum systems

### 04 Mar 2024 v1

@@ III Moments of inertia as the angular uncertainty measures
Now, let the body be a unit ring in the $X-Y$ plane, centered at the origin $O$. Assume further the ring to have a unit mass distributed along the ring with the angular mass density $p(\phi)$. From the definition (21) it then follows that the moment of inertia tensor of the ring about the origin $O$ is given by
@@
where ${\bf a}$ is the position vector of the point $G$ relative to the point $A$.
Moving back to the ring, we have ${\bf a}=(\langle C\rangle,\langle S\rangle,0)^{\rm T}$ and the theorem (23) implies
@@
where ${\bf n}_{xy}=(n_{x},n_{y})^{\rm T}$.
Similarity of the formula (27) with the variance $\langle(\Delta S_{\alpha})^{2}\rangle={\bf x}^{\rm T}{\bf\Gamma}{\bf x}$, Eq. (II), brings us to the idea to interpret all the angular uncertainty measures of the preceding section as moments of inertia of a ring about axes passing through its center of mass. The mechanical interpretation and the geometrical meaning of some of the measures is depicted in Figs. 1 and 2. Specifically, the dispersion (4) can be interpreted as a moment of inertia of the ring about an axis perpendicular to its plane, i.e., parallel with the unit vector ${\bf e}_{z}=(0,0,1)^{\rm T}$, and passing through its center of mass [29] (see Fig. 1 and the red axis in Fig. 2).
@@
This can be easily seen without any calculations in the frame rotated by an angle $-\mbox{arg}\langle E\rangle$ since the position of the center of mass is on the line $\langle S_{-\mbox{arg}\langle E\rangle}\rangle=0$ at the distance $\langle C_{-\mbox{arg}\langle E\rangle}\rangle=|\langle E\rangle|=\sqrt{1-D^{2}}$ from the origin (see Fig. 1). The moment of inertia of the ring with respect to the origin equals to 1 and the moment of inertia with respect to the center of mass is given by the parallel axis theorem as
@@
These are the special cases of moment of inertia with respect to the axis at the center of mass and oriented along the unit vector ${\bf n}=(\sin\theta\cos\Phi,\sin\theta\sin\Phi,\cos\theta)$ (see green axis in Fig. 2). A straightforward substitution into the formula (27) gives the result
@@
It is intriguing to note that variances used for quantification of uncertainties of position and momentum are nothing but moments of inertia of a probability distribution along the line (1D) with respect to perpendicular axis in the center of mass.
@@
How to understand all those results in the context of possible metrological applications? Moment of inertia is well defined quantity associated with the rotation of a rigid body. According to the inequalities (6) and (17) moments of inertia, namely $D^{2}\geq\gamma_{+}\geq{\bf m}^{\rm T}{\bf\Gamma}{\bf m}\geq{\bf x}_{\rm o}^{\rm T}{\bf\Gamma}{\bf x}_{\rm o}$ play the role of the ”uncertainties” of the angular variable. Obviously, dispersion as moment of inertia (28) seems to be the first choice but the Mathieu states are analytically intractable and cannot be fully exploited for further optimization. On the other hand measures constructed from higher order moments (17) and linked to moment of inertia (29) may seem to be forged ”artificially” but give rise to a simple tractable set of extremal von Mises states, which are effectively indistinguishable from Mathieu states. In the line with this interpretation the inequalities (6) and (17) cannot be seen as ”stronger” or ”weaker” since different measures are involved. However, both concepts, though ”slightly” different seem to be equivalent for all practical consequences and, all measures implied by inequalities (17) are mimicking dispersion when considering states close to optimal. Of course, differences may appear for ”non-optimal” states due to the differences in higher order moments.
It follows from this argument that states saturating the Robertson’s inequality are “minimum uncertainty states” if the concept of uncertainty is extended to projections of moment of inertia tensor with state-dependent orientation of axis of rotation: variance of sine (cosine) operator or dispersion are just extremal cases of more general formulae (30). Dispersion as a measure is in this sense exceptional since the axis is constant (perpendicular to $X-Y$ plane). Why is it worth to have some ”other” uncertainty relations? The answer is simple: They can characterise extremal states under different conditions and may provide tighter uncertainty relations! The tight form of generic uncertainty relations is formulated in Appendix B as an open problem, the solution of which is not an easy task. Here we will just stress pragmatical reasons why this could be of interest: If some particular physical platform of quantum rotor (see below) will allow to identify experimentally moments of angular variable $\langle E\rangle$ and $\langle E^{2}\rangle$, it could be valuable to find restrictions implied by quantum mechanics and to formulate stronger uncertainties. We will leave these issues for further research resorting to the simplest opportunity associated with the measure (29), which may find direct applications.
@@ IV Simultaneous detection of complementary variables
where $i,j=a,s$, and $\Delta A\equiv\sqrt{\langle(\Delta A)^{2}\rangle}$. The essence of the analysis is to find several branches which saturate the uncertainty relations and this is achieved by reordering the terms in the vector ${\bf d}_{ij}.$ As a result the uncertainty product can be assessed as
@@
where inequality $1$ follows from the Cauchy-Schwarz inequality, whereas inequality $2$ is a consequence of the uncertainty relation (6). The rationale behind the mathematical formulation stems from the fact that separate uncertainties $\langle(\Delta{\cal L})^{2}\rangle$ and $\mathcal{D}^{2}$ can be interpreted as norms of vectors, for which the Cauchy-Schwarz inequalities represent saturable bounds with several possible branches. The analysis hinges partially on numerical analysis due to the dependence on the state-dependent factor $\langle E_{a}\rangle.$ The results are summarized in Fig. 3. In Appendix C.1 we show further that the numerically found optimal value of the uncertainty product $\Pi_{\mathcal{D}}$ can be well approximated for ancilla prepared in the Mathieu state as ${\cal B}_{\mathcal{D}}$,
@@
where $D_{a,{\rm int}}^{2}\doteq 0.3$. The derivation was done here for dispersion and Mathieu extremal states but analogous formulation done for the uncertainty measure $\langle(\Delta\mathcal{S})^{2}\rangle$, Eq. (37), and von Mises states (see Fig. 4) shows the consistency of both the measures. In Appendix C.2 we show further that the optimal numerical bound is very well approximated by the function
@@
where $\tilde{D}_{a,{\rm int}}^{2}\doteq 0.167$ and the parameters of the signal and ancilla are matched by the condition $\kappa_{s}=\sqrt{I_{2}(2\kappa_{a})/I_{0}(2\kappa_{a})}\kappa_{a}$ [11]. This result has clear interpretation as condition matching the ”squeezing” of signal and ancillary systems. Notice in passing that for $x$ and $p$ quadrature operators of the harmonic oscillator both signal and ancillary systems are squeezed equally.
Disregarding the subtle differences between von Mises and Mathieu extremal states and optimal conditions for matching signal and ancillary systems, the ”almost optimal” simultaneous detection of complementary variables can be seen as detection of the von Mises signal state projected onto sequence of von Mises detector states (34) - a result fully analogous to the heterodyne detection for harmonic oscillator when squeezed state can be measured by projecting into squeezed (Gaussian) states. However the structure of ”optimal conditions” is much richer for quantum rotor with promises for future experimental realisations.
@@ V Quantum rotor in applications
There are several distinct and important physical platforms for implementing quantum rotor in the current technology. Detailed analysis of specific examples is beyond the scope of this contribution. Here we just point out commons features of all the examples, namely that uncertainties for complementary variables represent the first step towards optimal metrology. Note in passing the obvious example of systems with cylindrical symmetry, namely vortex beams, which was already discussed in Ref. [24], where the bounds for dispersion and variance of angular momentum were verified experimentally. It is intriguing to note that exponential angle operator can be constructed on 2 modes of transversal field associated with annihilation operators $a_{1},a_{2}$ as so called ”feasible” phase [32]
@@
As shown by Ban [33], there is an equivalent formulation of this problem in terms of relative number state representation. These concepts come from formal considerations of ”quantum phase operator” and provide exact (and intrinsically nonlinear) link between algebra of harmonic oscillator and quantum rotor.
Less obvious is the link between the problem of optimal shaping of the pulse in the time-frequency domain and the model of a quantum rotor. For the sake of simplicity we assume t-dependent signal on the $2\pi$ window in dimensionless coordinates, where $t$ plays the role of angular variable. Outside this interval let us fulfill for simplicity the periodicity condition $\psi(t+2\pi)=\psi(t).$ Such signals can be represented by discrete Fourier series in the selected time window
@@
If we define operators in t-representation as $L=-i\partial_{t},\quad E=e^{-it},$ we easily identify the commutation relation for $\mathfrak{e}$(2) algebra (3) with all the metrological consequences. In contrast with the ”standard” formulation where $x,p$ are related by Fourier transformation, the discrete $n$ and continuous t-variables are parameters in Fourier series. This problem will be addressed in a subsequent publication.
Contemporary quantum computing platforms for performing quantum operations are on circuits with Josephson junctions. Superconductors behave like macroscopic quantum mechanical systems. Only two quantities are required to describe the physics of a Josephson junction: the number imbalance of electrons $n$ (Cooper pairs) and $\theta$ is the relative phase between the two superconductors. The circuits built with superconducting components can carry currents without resistance because the carriers of the charge - electrons or holes near the Fermi energy level - are creating the Cooper pairs creating macroscopic coherent states as explained by BCS theory. Such a state can be described by complex-valued order parameter, the phase of which is essential for the physics of super-conducting qubits [34]. The standard explanation in solid state physics relies on the canonical pair "$[\hat{\theta},\hat{n}]"=i,$ though it is known that this form is mathematically not rigorous due to the periodicity of angular variable [35, 14, 36]. The quotation marks are indicating the potential problems: the number of Cooper pairs $n$ tunneling through the barrier in Josephson junction can be negative while $\theta$ linked with gauge phase over the barrier is periodic, bearing the striking similarity with ”phase-number operators”. However, the relation $e^{i\theta}ne^{-i\theta}=n-1$ is free of those problems but this is nothing else but the algebraic expression for Euclidean algebra $\mathfrak{e}$(2), Eq. ( 3), for the exponential operator $E=e^{-i\theta}.$ But there are still other analogies linking the uncertainty relations for the quantum rotor. Under the conditions of super-conductivity only Cooper pairs tunnel coherently in the superconducting junction, and the system is described by the Hamiltonian of island -base qubit with capacitance coupled to Josephson junction [34] called also Cooper pair box
@@
where $n$ is number operator for excess Cooper pair charges in the island, $\theta$ is the phase of the superconducting order parameter of the island, $E_{C}$ is the charging energy and $E_{J}$ is the Josephson junction energy. As pointed out in [35], the Cooper box is to quantum circuit physics what the hydrogen atom is to atomic physics. But this Hamiltonian of a quantum rotor [13] is the same as the extremal equation (5) for states minimising uncertainty for the variance of angular momentum and dispersion (6). In other words, projections into the (displaced) ground state of the quantum rotor have the same meaning for quantum tomography of super-conduction qubits as projections into (displaced) ground state of harmonic oscillator, i.e., projections into squeezed states of electromagnetic field. Since Mathieu states can be very closely approximated by von Mises states the theory presented here provides deep analogy between quantum metrology of harmonic oscillator and quantum rotor. This is facilitated by the generalised measurement of commuting pair (33), where the signal and ancillary loops should be coupled to more complex circuits allowing to detect super-imposed (or subtracted) complementary variables from both loops. A practical goal of the results presented here could be the design of optimised detection and diagnostics of super-conducting qubits such as charge-, phase- or transmon qubit states [37]. In order to reach such an ambitious goal concentrated effort is surely needed utilising the potential advantages of both quantum harmonic oscillator and rotor.
@@ VI Conclusion
The quantum rotor - the second simplest model of complementary variables after harmonic oscillator is of paramount yet under-appreciated importance in contemporary quantum science. Though rotational symmetry is a well understood concept, insight into the canonical pair of angular momentum and angular variable was obscured by problems stemming from inadequate description of complementary variables and the lack of saturable uncertainty relations. We show that moment of inertia with the probability distributed on the unit ring with respect to various rotation axes going through center of mass can be used for meaningful quantification of angular variable and for the formulation of variations of the uncertainty principle. This unifying approach allows to formulate the metrology of quantum rotor along the lines of those known for harmonic oscillator. Particular attention is payed to relations saturated by Mathieu or von Mises states allowing to formulate concept of optimal detection of complementary variables for system like pulses or super-conducting qubits with Josephson junction. Quantum information processing adopting both basic models of harmonic oscillator and quantum rotor may bring new and challenging ideas for processing of quantum information due to the inherent nonlinearities between both models: what is linear operation for harmonic oscillator is nonlinear for rotor and vice versa.
@@ Acknowledgements.
Appendix A Derivation of moment of inertia for arbitrary axis
Here we give a simple and straightforward proof of the moment of inertia (30) without using tensor calculations. Assume the coordinate system of the unit ring, where $X$ axis goes through its center of mass, vectors of axis of rotation ${\bf n}=(\cos\Phi\sin\theta,\sin\Phi\sin\theta,\cos\theta)$ and position on the unit ring at the $X-Y$ plane ${\bf e}=(\cos\varphi,\sin\varphi,0).$ The moment of inertia of the ring with respect to the axis at origin is given as
$M_{0}=\langle\sin^{2}\vartheta\rangle_{\varphi},$ (49)

### 13 Mar 2024 v2

@@ III Moment of inertia as the angular uncertainty measure
Let the body be a unit ring in the $X-Y$ plane, centered at the origin $O$ with the unit mass distributed along the ring with the angular mass density $p(\phi)$. From the definition (21) it follows that the moment of inertia tensor of the ring about the origin $O$ is given by
@@
where ${\bf a}$ is the position vector of the point $G$ relative to the point $A$. For the ring we have ${\bf a}=(\langle C\rangle,\langle S\rangle,0)^{\rm T}$ and the theorem (23) implies
@@
where ${\bf n}_{xy}=(n_{x},n_{y})^{\rm T}$. Due to the formula (27), the variance $\langle(\Delta S_{\alpha})^{2}\rangle={\bf x}^{\rm T}{\bf\Gamma}{\bf x}$, Eq. (II), can be interpreted as the moment of inertia of a ring about axes passing through its center of mass. The mechanical interpretation and the geometrical meaning of some of the measures is depicted in Figs. 1 and 2. Specifically, the dispersion (4) is the moment of inertia of the ring about an axis perpendicular to its plane, i.e., parallel with the unit vector ${\bf e}_{z}=(0,0,1)^{\rm T}$, and passing through its center of mass [31] (see Fig. 1 and the red axis in Fig. 2)
@@
The measures discussed in previous section are the special cases of moment of inertia with respect to the axis at the center of mass and oriented along the unit vector ${\bf n}=(\sin\theta\cos\Phi,\sin\theta\sin\Phi,\cos\theta)$ (see green axis in Fig. 2). A straightforward substitution into the formula (27) gives the result
@@
Elementary derivation of generic moment of inertia is given in Appendix A. It is also intriguing to note that standard variances used for quantification of uncertainties of position and momentum are nothing but moments of inertia of a probability distribution along the line (1D) with respect to perpendicular axis in the center of mass.
@@
How to understand all those results in the context of possible metrological applications? Moment of inertia is well defined quantity associated with the rotation of a rigid body. According to the inequalities (6) and (17) moments of inertia, namely $D^{2}\geq\gamma_{+}\geq{\bf m}^{\rm T}{\bf\Gamma}{\bf m}\geq{\bf x}_{\rm o}^{\rm T}{\bf\Gamma}{\bf x}_{\rm o}$ play the role of the uncertainties of the angular variable. Obviously, dispersion as moment of inertia (28) seems to be the first choice but the Mathieu states are analytically intractable and cannot be fully exploited for further optimization. On the other hand measures constructed from higher order moments (17) and linked to moment of inertia (29) may seem to be forged artificially but give rise to a simple tractable set of extremal von Mises states, which are effectively indistinguishable from Mathieu states. In the line with this interpretation the inequalities (6) and (17) cannot be seen as stronger or weaker since different measures are involved. However, both concepts, though slightly different seem to be equivalent for all practical consequences and, all measures implied by inequalities (17) are mimicking dispersion when considering states close to optimal. Of course, differences may appear for non-optimal states due to the differences in higher order moments.
It follows from this argument that states saturating the Robertson’s inequality are “minimum uncertainty states” if the concept of uncertainty is extended to projections of moment of inertia tensor with state-dependent orientation of axis of rotation: variance of sine (cosine) operator or dispersion are just extremal cases of more general formulae (30). Dispersion is in this sense exceptional since the axis is constant (perpendicular to $X-Y$ plane). Why is it worth to have some other uncertainty relations? The answer is simple: They can characterise extremal states under different conditions and may provide tighter uncertainty relations! The tight form of generic uncertainty relations is formulated in Appendix B. The extremal states are attributed to solutions of Hill equation [27], which is a generalization of Mathieu equation. Here we will just stress the reasons why this could be of interest: If some particular physical platform of quantum rotor (see examples below) will allow to identify experimentally moments of angular variable $\langle E\rangle$ and $\langle E^{2}\rangle$, it could be valuable to find restrictions implied by quantum mechanics and to formulate stronger uncertainties. We will leave these issues for further research resorting to the simplest opportunity associated with the measure (29), which may find direct applications now.
@@ IV Simultaneous detection of complementary variables
where $i,j=a,s$, and $\Delta A\equiv\sqrt{\langle(\Delta A)^{2}\rangle}$. The rationale behind the definition of ${\bf d}_{ij}$ stems from the fact that separate uncertainties $\langle(\Delta{\cal L})^{2}\rangle$ and $\mathcal{D}^{2}$ can be interpreted as norms of vectors, for which the Cauchy-Schwarz inequalities will represent saturable bounds with several possible branches. As a result the uncertainty product can be assessed as
@@
where inequality $1$ follows from the Cauchy-Schwarz inequality, whereas inequality $2$ is a consequence of the uncertainty relation (6). The analysis hinges partially on numerical analysis due to the dependence on the state-dependent factor $\langle E_{a}\rangle.$ The results are summarized in Fig. 3. In Appendix C.1 we show that the numerically found optimal value of the uncertainty product $\Pi_{\mathcal{D}}$ can be well approximated for ancilla prepared in the Mathieu state as ${\cal B}_{\mathcal{D}}$,
@@
where $D_{a,{\rm int}}^{2}\doteq 0.3$. The derivation was done here for dispersion and Mathieu extremal states but analogous formulation done for the uncertainty measure $\langle(\Delta\mathcal{S})^{2}\rangle$, Eq. (37), and von Mises states (see Fig. 4) shows the consistency of both the measures. In Appendix C.2 we further show that the optimal numerical bound is very well approximated by the function
@@
where $\tilde{D}_{a,{\rm int}}^{2}\doteq 0.167$ and the parameters of the signal and ancilla are matched by the condition $\kappa_{s}=\sqrt{I_{2}(2\kappa_{a})/I_{0}(2\kappa_{a})}\kappa_{a}$ [11]. This result has clear interpretation as condition matching the squeezing of signal and ancillary systems. Notice that for $x$ and $p$ quadrature operators of the harmonic oscillator both signal and ancillary systems are squeezed equally.
Disregarding the subtle differences between von Mises and Mathieu extremal states and optimal conditions for matching signal and ancillary systems, the ”almost optimal” simultaneous detection of complementary variables can be seen as detection of the von Mises signal state projected onto sequence of von Mises detector states (34) - a result fully analogous to the heterodyne detection for harmonic oscillator when squeezed state can be measured by projecting into squeezed (Gaussian) states. However the structure of optimal conditions is much richer for quantum rotor with promises for future experimental realisations.
@@ V Quantum rotor in applications
There are several distinct and important physical platforms for implementing quantum rotor in the current technology. Detailed analysis of specific examples is beyond the scope of this contribution. Here we just point out commons features of all the examples, namely that uncertainties for complementary variables represent the first step towards optimal metrology.
As the first example let us mention system with cylindrical symmetry-vortex beams. Uncertainties for dispersion and variance of angular momentum were analyzed theoretically and verified experimentally in Ref. [26]. It is intriguing to note that exponential angle operator can be constructed on two modes of transversal field associated with annihilation operators $a_{1},a_{2}$ as so called ”feasible” phase [34]
@@
As shown by Ban [35], there is an equivalent formulation of this problem in terms of relative number state representation. These concepts come from formal considerations of quantum phase operator and provide exact (and intrinsically nonlinear) link between algebra of harmonic oscillator and quantum rotor.
Less obvious is the link between the problem of optimal shaping of the pulse in the time-frequency domain within the model of quantum rotor. For the sake of simplicity we assume t-dependent signal on the $2\pi$ window in dimensionless coordinates, where $t$ plays the role of angular variable. Outside this interval let us fulfill the periodicity condition $\psi(t+2\pi)=\psi(t).$ Such signals can be represented by discrete Fourier series in the selected time window
@@
If we define operators in t-representation as $L=-i\partial_{t},\quad E=e^{-it},$ we easily identify the commutation relation for $\mathfrak{e}$(2) algebra (3) with all the metrological consequences. In contrast with the standard formulation where $x,p$ are related by Fourier transformation, the discrete $n$ and continuous t-variables are parameters in Fourier series. This problem will be addressed in a subsequent publication.
Contemporary quantum computing platforms for performing quantum operations are based on circuits with Josephson junctions [36, 37]. Superconductors behave like macroscopic quantum mechanical systems. Only two quantities are required to describe the physics of a Josephson junction: the number imbalance of electrons $n$ (Cooper pairs) and $\theta$ is the relative phase between the two superconductors. The circuits built with superconducting components can carry currents without resistance because the carriers of the charge - electrons or holes near the Fermi energy level - are creating the Cooper pairs behaving like macroscopic coherent states as explained by BCS theory. Such a state can be described by complex-valued order parameter, the phase of which is essential for the physics of super-conducting qubits [38, 39]. The standard explanation in solid state physics relies on the canonical pair "$[\hat{\theta},\hat{n}]"=i,$ though it is known that this form is mathematically not rigorous due to the periodicity of angular variable [40, 16, 41]. The quotation marks are indicating the potential problems: the number of Cooper pairs $n$ tunneling through the barrier in Josephson junction can be negative while $\theta$ linked with gauge phase over the barrier is periodic, bearing the striking similarity with phase-number operators. However, the relation $e^{i\theta}ne^{-i\theta}=n-1$ is free of those problems but this is nothing else but the algebraic expression for Euclidean algebra $\mathfrak{e}$(2), Eq. (3), for the exponential operator $E=e^{-i\theta}.$
There are still other analogies with uncertainty relations for the quantum rotor. Under the conditions of super-conductivity only Cooper pairs tunnel coherently in the superconducting junction, and the system is described by the Hamiltonian of island-base qubit with capacitance coupled to Josephson junction [38] called also Cooper pair box
@@
where $E_{C}$ is the charging energy and $E_{J}$ is the Josephson junction energy. As pointed out in [40], the Cooper box is to quantum circuit physics what the hydrogen atom is to atomic physics. But this Hamiltonian of a quantum rotor [15] is the same as the extremal equation (5) for states minimising uncertainty for the variance of angular momentum and dispersion (6). In other words, projections into the (displaced) ground state of the quantum rotor have the same meaning for quantum tomography of super-conducting qubits as projections into (displaced) ground state of harmonic oscillator, i.e., projections into squeezed states of electromagnetic field. Since Mathieu states can be very closely approximated by von Mises states the theory presented here provides deep analogy between quantum metrology of harmonic oscillator and quantum rotor. This is facilitated by the generalised measurement of commuting pair (33), where the signal and ancillary loops should be coupled to more complex circuits allowing to detect super-imposed (or subtracted) complementary variables from both loops in similar way as it is done in quantum optics with homodyne or heterodyne detection. Practical and ambitious goal of theoretical analysis presented in this research could be the design of optimised detection and diagnostic schemes for circuits built on the basis of Josephson junction, which provides platforms for charge-, phase- or fluxons super-conducting qubits [42, 39].
@@ VI Conclusion
Uncertainty relations for quantum rotor can be formulated by means of moment of inertia of the probability distributed along the unit ring with respect to various rotation axes going through center of mass. Unifying approach allows to develop quantum metrology at ultimate quantum limits analogously to results known for harmonic oscillator. Particularly, the role of squeezed states is played by von Mises states allowing to formulate optimal detection of complementary variables for system described naturally by generators of the Euclidean group E(2) like vortex beams, pulses in time-frequency domain or super-conducting qubits with Josephson junctions. Quantum information processing may benefit from both fundamental models of harmonic oscillator and quantum rotor due to their intrinsic non-linear relationship.
@@ Acknowledgements.
Appendix A Elementary derivation of moment of inertia for arbitrary axis
Here we give elementary and straightforward derivation of the moment of inertia (30) without using tensor calculations. Assume the coordinate system of the unit ring, where $X$ axis goes through its center of mass, vectors of axis of rotation ${\bf n}=(\cos\Phi\sin\theta,\sin\Phi\sin\theta,\cos\theta)$ and position on the unit ring at the $X-Y$ plane ${\bf e}=(\cos\varphi,\sin\varphi,0).$ The moment of inertia of the ring with respect to the axis at origin is given as
$M_{0}=\langle\sin^{2}(\vartheta)\rangle_{\varphi},$ (49)